# RepLiQA: A Question-Answering Dataset for Benchmarking LLMs on Unseen Reference Content

### v1

@@ 4 Benchmarking LLMs with RepLiQA on Reading Comprehension
We select eighteen widely-used LLMs: GPT-3.5 and GPT-4o by OpenAI (Achiam et al., 2023), Llama 3 by Meta (Touvron et al., 2023; AI@Meta, 2024) with both 8 and 70 billion parameters, Gemini 1.0 and 1.5 by Google (Gemini Team, 2023), two variants of WizardLM by Microsoft (Xu et al., 2024a; Luo et al., 2024), Mistral and Mixtral variants by MistralAI (Jiang et al., 2023), Command R and Command R+ by Cohere (2024), Arctic by Snowflake (2024), and  Anthropic (2024)’s Claude Haiku and Sonnet. Inference is carried out using OpenRouter, [^5] https://openrouter.ai/ which serves as a unified framework enabling access to multiple LLM providers through a single API. Our entire benchmarking evaluation has cost approximately USD 5k.
@@
For question answering, we follow the evaluation protocol by Adlakha et al. (2023) and measure performance metrics such as the F1 score and recall. Around 20% of questions in RepLiQA are not answerable from provided documents, in which situation we expect (and prompt) models to reply with unanswerable. We further evaluate models in terms of their ability to detect such situations and abstain from answering. For topic retrieval, models were prompted to determine the topic or document category out of the occurrences in the dataset (cf. Figure 3), with the set of candidates passed through the prompt. Further evaluation details, such as the prompts we used, can be found in Appendix F.
@@ 4.1 Question-Answering
In Figure 4 we report the question answering results on RepLiQA0 and investigate how they differ from those on TriviaQA. We observe that all models perform significantly better on TriviaQA than on RepLiQA, hinting to the fact that models might rely on memory and not acquired reading skills to solve these tasks. Recall gaps, shown in the bottom-left plot of Figure 4, underline the extent to which model performance drops in the situation where knowledge obtained during training is not useful (e.g., Command R+’s recall is halved). For some models the gap is not as dramatic: Claude Sonnet, Llama 3 8B and Arctic behave similarly across the two datasets, although their performance are suboptimal. The best-performing model on RepLiQA is Mistral Large whose recall gap is also amongst the smallest.
@@
Figure 5: Impact of the presence or absence of context when answering questions, measured using recall, for various models on both RepLiQA and TriviaQA. The results on RepLiQA are restricted to the questions whose answers are not unanswerable.
@@ 4.2 Effect of Scaling Model Size
4.3 Testing Abstention Ability of Models
Figure 7: Comparison of selected models on the unanswerable detection task (higher is better) on RepLiQA, with or without context. When including the context, scores are computed only for unanswerable questions. A perfect model would score 100%.
@@ 4.3 Testing Abstention Ability of Models
Interestingly, all models tend to refrain from answering unanswerable questions more often when context is provided than when one is not provided, and they tend to come up with an answer, even if it is not possible to do so. We highlight that, as shown in Appendix F, prompts include instructions for models not to answer in situations where documents (or knowledge obtained during pre-training) do not offer useful information.

### v2

@@ 4 Benchmarking LLMs with RepLiQA on Reading Comprehension
We select eighteen widely-used LLMs: GPT-3.5 and GPT-4o by OpenAI (Achiam et al., 2023), Llama 3 by Meta (Touvron et al., 2023; AI@Meta, 2024) with both 8 and 70 billion parameters, Gemini 1.0 and 1.5 by Google (Gemini Team, 2023), two variants of WizardLM by Microsoft (Xu et al., 2024a; Luo et al., 2024), Mistral and Mixtral variants by MistralAI (Jiang et al., 2023), Command R and Command R+ by Cohere (2024), Arctic by Snowflake (2024), and  Anthropic (2024)’s Claude Haiku and Sonnet.
Inference is carried out using OpenRouter, [^5] https://openrouter.ai/ which serves as a unified framework enabling access to multiple LLM providers through a single API. Our entire benchmarking evaluation has cost approximately USD 5k.
@@
For question answering, we follow the evaluation protocol by Adlakha et al. (2023) and measure performance metrics such as the F1 score and recall. Around 20% of questions in RepLiQA are not answerable from provided documents, in which situation we expect (and prompt) models to reply with unanswerable. We further evaluate models in terms of their ability to detect such situations and refuse to reply.
For topic retrieval, models were prompted to determine the topic or document category out of the occurrences in the dataset (cf. Figure 3), with the set of candidates passed through the prompt. Further evaluation details, such as the prompts we used, can be found in Appendix F.
@@ 4.1 Question-Answering
Figure 5: Side-by-side performances for each model on RepLiQA and TriviaQA.
In Figure 4 we report the question answering results on RepLiQA0 and investigate how they differ from those on TriviaQA. We observe that all models perform significantly better on TriviaQA than on RepLiQA, hinting to the fact that models might rely on memory and not acquired reading skills to solve these tasks. Note that standard errors for recall are too small to be readable in the figures (0.25% to 0.28% for RepLiQA and 0.29% to 0.50% for TriviaQA). Recall gaps, shown in the bottom-left plot of Figure 4, underline the extent to which model performance drops in the situation where knowledge obtained during training is not useful (e.g., Command R+’s recall is halved). For some models the gap is not as dramatic: Claude Sonnet, Llama 3 8B and Arctic behave similarly across the two datasets, although their performance is suboptimal. The best-performing model on RepLiQA is Mistral Large whose recall gap is also amongst the smallest. We show further evidence of performance differences across the two datasets in Figure 5. For all models we evaluated across the situations where reference documents are based on information available during training or not. Moreover, in Figure 6, we show the performance distributions across the two datasets. We observe a pronounced skewness towards high recall on TriviaQA. Extra results with a models specialized on Question-Answering are reported in Appendix I.
Figure 6: Violin plots depicting differences in performance distributions across RepLiQA and TriviaQA. We observe a pronounced skewness towards high recall on TriviaQA.
@@
Figure 7: Impact of the presence or absence of context when answering questions, measured using recall, for various models on both RepLiQA and TriviaQA. The results on RepLiQA are restricted to the questions whose answers are not unanswerable. Note that recall can be non-zero for a model that only answers wrongly if it outputs a few tokens that appear in the ground-truth answers.
@@ 4.2 Effect of Scaling Model Size
4.3 Testing The Ability of Models to Admit Lack of Knowledge
@@
Figure 9: Comparison of selected models on the unanswerable detection task (higher is better) on RepLiQA, with or without context. When including the context, scores are computed only for unanswerable questions. A perfect model would reach 100%.
Interestingly, all models tend to refuse answering unanswerable questions more often when context is provided than when one is not provided, and they tend to come up with an answer, even if it is not possible to do so. We highlight that, as shown in Appendix F, prompts include instructions for models to refuse to answer in situations where documents (or knowledge obtained during pre-training) do not offer useful information.
@@ 5 Conclusion
We also highlight a possible bias toward Indian English with respect to other variants of English within RepLiQA due to the demographics of annotators. That being said, this kind of bias (if present) would be more important in a dataset meant for training, which RepLiQA is not. RepLiQA is meant for benchmarking a model’s ability to answer (or identify as unanswerable) questions based on a never-seen-before context. A priori, this capability should be orthogonal to the model’s performances in different English dialects. Moreover, our manual inspection of the dataset indicates that a rather “international” English dialect is used.
@@ Appendix G Related Work
Appendix H Topic classification results
We tested to what extent topics in RepLiQA are well defined in the sense that there’s little confusion on a topic given a document. To do so, we trained a topic classifier given by a pre-trained LongFormer encoder [Beltagy et al., 2020]. Trained classifiers achieved test accuracy greater than 95%, which suggests that overlap is not a reason for concern. Code implementing such experiment can be found at: https://github.com/ServiceNow/repliqa/blob/main/repliqa_topic_classifier.ipynb. In Figure 12, we report a confusion matrix showing little topic ambiguity.
Figure 12: Confusion matrix for predictions of a topic classifier trained on RepLiQA.
Appendix I Evaluating models fine-tuned for question-answer tasks
While our main evaluations focused on general purpose language models, we additionally evaluated a model that was fine-tuned specifically for question-answering tasks. Namely, we evaluated Llama3-ChatQA-1.5-8B (indicated as fine-Tuned in the bottom row of table 8). We observed a performance degradation on RepLiQA and a slight improvement on TriviaQA, which is in alignment with evaluations of other fine-tuned models on a pre-release version of RepLiQA and other datasets, as reported in [Monteiro et al., 2024]. Note that we closely followed the prompt format recommended in https://huggingface.co/nvidia/Llama3-ChatQA-1.5-8B. Inference code can be found at: https://github.com/ServiceNow/repliqa/blob/main/hf_qa_eval.ipynb.
Table 8: Question-Answer performance of a fine-tuned model in terms of recall.
RepLiQA TriviaQA
Claude 3 Sonnet 66.54 75.41
Mistral Large 72.29 86.79
LLama 3 8B 54.82 70.95
LLama 3 8B (Fine Tuned) 57.19 63.51